A and B are playing tag . A stay still while B runs from the same spot at a rate 2 meters per second. After 60 seconds B keeps running . A now runs after him at a rate of 3 meters per second. How long has A run when he caught up to B
My solution
3 (t-60) = 2t
3t - 180 =2t
3t-2t =180
t = 180

Answer by lwsshak3(11628)

A and B are playing tag . A stay still while B runs from the same spot at a rate 2 meters per second. After 60 seconds B keeps running . A now runs after him at a rate of 3 meters per second. How long has A run when he caught up to B
My solution
3 (t-60) = 2t
3t - 180 =2t
3t-2t =180
t = 180
***
After 60 seconds, B has covered a distance=2*60=120 m
3-2=1 m/sec=rate at which A is catching up to B
A's travel time=120/1=120 seconds
How long has A run when he caught up to B? 120 seconds